OpenStudy (anonymous):

find g(1)+g(2)+g(3)...... + g(51) g(x)= 5-x

OpenStudy (kirbykirby):

g(1) = 5-1=4 g(2) = 5-2=3 g(3) = 5-3=2 ... g(51) = 5 - 51 = -46 So you are essentially finding the sum: 4 + 3 + 2 + 1 + 0 - 1 -2 ... - 46

OpenStudy (anonymous):

well, i understand that, i was mainly wondering if there was a formula for doing that?

OpenStudy (kirbykirby):

You can also just use summation notation on the formula and do \[\large \sum_{x=1}^{51}(5-x)=\sum_{x=1}^{51}5-\sum_{x=1}^{51}x \] And use the formula: \[\large \sum_{x=1}^{n}x=\frac{n(n+1)}{2} \]

OpenStudy (anonymous):

ok, so it would be 51(52)/2 ?

OpenStudy (kirbykirby):

yes, for the second summation

OpenStudy (anonymous):

the first it would be is 5(6)/2

OpenStudy (kirbykirby):

no not quite.. the formula works if you have the variable "x" in front of the summation, the 5 you have is a constant. When dealing with constants, essentially the formula is telling you you are adding "5" 51-times.

OpenStudy (kirbykirby):

essentially you need the variable you are summing to match with the index of summation: \[\Large\sum_{\color{red}{x=1}}^{51}\color{red}{x} \], the variables agree, so you can use the above formula. But if you had: \[\Large\sum_{\color{red}{x=1}}^{51}\color{blue}{a} \], the variables don't agree, so the "a" is constant with respect to the summation index. So in that case you calculate the sum as \(51\cdot a\)

OpenStudy (anonymous):

ok, so how would i do that?

OpenStudy (kirbykirby):

in your case the "a" is just 5, so you find the first summation as \(51(5)\)

OpenStudy (anonymous):

oh ok i figured out how to do it on my calculator, thank you for your help though'

OpenStudy (kirbykirby):

:)